### v1

@@ 4.2 Results and Analysis
GPT4-AllRules is the best performing model on the Dimension questions, although its average accuracy (0.533) is marginally better than the naive baseline (random yes-no guessing). We noticed several recurring issues that caused the model to not answer questions correctly. First, the model sometimes struggled to reference the relevant rule (as seen in the Retrieval questions), and this meant that the model would not be set-up to answer the question correctly. For example, when asked three different questions pertaining to Rule F.5.7.5, GPT4-AllRules quotes the rule differently in each of its three explanations: one time it quotes the correct rule, another time it mistakenly quotes the text of the preceding rule (F.5.7.4), and the third time it mistakenly quotes the text of a rule two rules ahead (F.5.7.7). Second, we notice that GPT4-AllRules sometimes has difficulty in extracting dimensions from the engineering drawing. For example, in two scale bar questions each which has a scale bar of 3202.4 mm, the model says that the scale bar is 1000 mm in one instance and 500 mm in another instance. Third, we note that the model faces difficulty in answering questions where the drawing contains two dimensions, and the correct answer can only be obtained by adding or subtracting the two together. For example, in two questions about the same rule, the model scrapes one of two dimensions from the provided engineering drawing and uses it to answer the compliance question, when the difference between the two dimensions should actually be used.
The drop in accuracy score between GPT4-AllRules and GPT4-RAG can largely be attributed to the fact that the RAG does not provide the model with the relevant portion of the rule document (as seen in the Retrieval questions). In many cases when this happens, GPT4-RAG refuses to provide a yes/no answer (as in the example in Figure 3), resulting in an accuracy score of zero for that question. LLaVA-RAG receives the same portions of the rules as GPT4-RAG, but it does not indicate when the portion of the rules doesn’t contain the rule-in-question. As a result, LLaVA5-RAG provides yes/no answers to most questions (51/60) while GPT4-RAG provides yes/no answers to many fewer questions (21/60). Because providing a yes/no answer improves the chances of guessing correctly (while refusing to answer automatically results in a score of zero), LLaVA-RAG receives an inflated average accuracy score on the dimension questions compared to GPT4-RAG. We also observe that many of the explanations that the models provide are significantly longer than the reference explanations we have (see Figure 3). It is therefore difficult to capture the correctness of explanations through the Bleu and Rogue scores. The interpretability of these scores would be improved by obtaining more human reference explanations, so as to better capture the distribution of possible explanations that could be considered correct.
As explained in Section 3.1.2, one-third of the questions comprising the Dimension QAs used scale-bars in the engineering drawings, while the other two-thirds used direct dimensions. We were curious to see what impact these two different dimensioning systems had on model performance. As seen in Table 4, all models perform better on the direct-dimensioned drawings than on the scale-bar drawings. The models seemed to struggle with the scale-bar, sometimes indicating in their explanations that they were using it for ‘‘rough"or ‘‘estimated"dimensions rather than precise ones and in some cases explaining that their ‘‘image capabilities…do not include measuring dimensions."As explained in Section 3.1.4, half of the Dimension questions were given additional context that we believed would help in answering the question while the other half were not. Surprisingly, we did not see any obvious trends in model performance with versus without the additional context. GPT4-AllRules and LLaVA-RAG performed worse with additional context, while GPT4-RAG performed better. More investigation into the effect of additional context is needed.
Таблица 4: Effect of dimensioning system used in engineering drawings on model accuracy on Dimension QAs. These results highlight the fact that the MLLMs tested are better able to answer questions about engineering drawings that are direct-dimensioned rather than those that have a scale-bar.
ACC by Dimension System GPT4-AllRules GPT4-RAG LLaVA-RAG
Direct 0.663 0.45 0.41
Multi-mention 0.275 0 0.400
@@
All of the MLLMs tested received the same accuracy score on the Functional Performance questions, although they did not provide identical yes/no responses for each question. Similar to the Dimension questions, we notice that the models sometimes cite the wrong rule, setting up the rest of the question for failure. We also notice some instances of not extracting information correctly from the provided images (e.g. not identifying that an image of FEA has a numerical scale bar or not correctly identifying the peak force in a force-displacement graph). We also observe some instances where a model makes additional assumptions about the question that are not incorrect, but encourages it to choose the wrong yes/no answer. For example, in one question, GPT4-AllRules is presented with a graph showing gearbox temperature over time, which approaches but does not exceed 60 °$\degree$°C. The rule in question specifies that the maximum temperature should not exceed 60 °$\degree$°C; while data provided does not, the model explains that it could if tested for more time, so it is not rule compliant. This is a valid answer, but not what we intended. Additional context in the prompt could help improve the clarity of these questions. Moreover, this set of questions was limited to just 16 due to limited resources and the time and expertise required to obtain and write these questions. Future work will focus on expanding this subset to better characterize model capabilities on more functional performance-type questions.
Recommendations for Models for This Benchmark
@@
In the results presented thus far, we have demonstrated how two state-of-the-art models, with and without simple RAG, perform against our benchmark. Here, we present some observations on how one might modify these models to achieve improved performance on our benchmark.
First, experimentation with the RAG model would be worthwhile so that relevant portions of the rule document are better retrieved. While we generally saw better performance when the full rule document was included in the model’s context window instead of provided to the model via RAG (GPT4-AllRules vs. GPT4-RAG), including the full rule document in the context of every question is very expensive (around \$800 for all questions in the dataset, at the time of writing, when using the latest vision model through the OpenAI API). As such, effective RAG could help reduce the computational burden. As seen in the results presented for the Retrieval questions, the simple LlamaIndex RAG used in this paper often does not provide the relevant section of the rule document to the model. Since the benchmark contains VQAs, we expect that models coupled with multimodal RAG – where the QA image could help in selecting the relevant portion of the rule document – could help improve scores on the benchmark. The RAG could also perhaps be improved by experimenting with different chunking methods. Nevertheless, as models become cheaper, more computationally efficient, and trained with larger context windows, RAG approaches might become less useful.
Second, improving the quality of the instructions provided to the MLLM models, by using techniques such as prompt-engineering, might result in improved performance on the benchmark. For example, extracting specific portions of the prompt to be fed to the RAG would likely improve the scores of models with RAG. Few-shot learning or including chain of thought reasoning [22] may also improve model scores. Third, we suspect that fine-tuned models would exhibit improved performance on the benchmark. Since many tasks in the benchmark require that the model can first extract a relevant rule and then answer a question, fine-tuning a model on the Retrieval QAs may result in improved performance across the entire benchmark. This capability could also be obtained by hybrid models that combine rule-based search and deep learning. Lastly, while we presented the results of two base models in this work, many other closed and open-source models could be evaluated on our benchmark and might perform better.
5 FUTURE WORK AND LIMITATIONS
@@
Engineering design, inherently grounded in practical applications, necessitates a deep understanding of technical documents to ensure that designs not only meet but also adhere to stringent requirements and standards. One of the critical insights emerging from our benchmark evaluations is the realization that modern AI models are still in the nascent stages of truly understanding engineering documentation. This limitation highlights a significant gap in the path toward fully automated design processes. As highlighted in the prior section, there is therefore substantial future work to be done to improve MLLM models’ abilities to solve engineering design and design requirement problems.
While this benchmark presents a first step into formally evaluating a model’s ability to interpret and understand design requirement documentation, the size of our current benchmark is limited to just one requirement document, and the creation costs (requiring expert manual validation) pose limitations on the generalizability and scalability of the dataset. Moreover, the metrics used to evaluate models against the benchmark could be further improved, as Bleu and Rouge have known limitations and qualitatively do not perfectly capture the correctness of the generated answers.

### v2

@@ 4.2 Results and Analysis
Of the models tested, GPT-4o-AllRules performs the best in terms of accuracy on the Dimension questions. It performs well, with an accuracy score of 0.825. We looked closer at where GPT-4o-AllRules struggled in answering the basic (no additional context, see Section 3.1.4) Dimension questions, to gain insight into its failure modes. For five questions that it answered incorrectly, GPT-4o-AllRules struggled to compute accurate dimensions using a scale bar. Models’ understanding of scale bar versus directly dimensioned engineering drawings is explored further in Section 5. Two of the questions that GPT-4o-AllRules answered incorrectly featured engineering drawings that displayed two dimensions; the correct answer to the question required that the model take a sum or difference between the two dimensions. The model struggled on these questions because it scraped a single dimension as the answer, rather than scraping both dimensions and combining them to get the answer. In another two questions GPT-4o-AllRules answered incorrectly, it mistakenly identified which vehicle components the dimensions were spanning in the engineering drawings. We also observed an instance of the model extracting an incorrect number from the engineering drawing image (9 instead of 0.9) and an instance of the model referencing the wrong rule from the rule document. These findings indicate that MLLMs’ still have room for improvement when it comes to skills – like visual component recognition, OCR, and dimensional computation – that are needed to answer questions pertaining to engineering drawings.
Of the models tested, GPT-4o-AllRules has the highest Similarity score for its Dimension explanations, matching its best-performer accuracy score. However, it is not the best-performer on the BLEU and ROUGE metrics, where Gemini-1.0-RAG has the higher (or tied, in the case of ROUGE) scores. We observe that many of the MLLMs’ explanations are significantly longer than the reference explanations we collected (see Figure 3). We suspect that Gemini-1.0-RAG’s high scores on the BLEU and ROUGE metrics are because its explanations tended to be shorter, more comparable in length to the human-generated, reference explanations. It is therefore challenging to capture the correctness of explanations through the BLEU and ROUGE scores. The interpretability of these scores would be improved by obtaining more human reference explanations, so as to better capture the distribution of possible explanations that could be considered correct.
@@
GPT-4o-AllRules once again performed the best on this subset of the benchmark, of the models tested. Its accuracy score was quite high, answering only one of the 16 questions incorrectly. The question the model answered incorrectly pertains to an anthropomorphic data chart showing requirements for human size (height, weight, etc.) for the vehicle and another chart that shows the human sizes the built vehicle can actually accommodate. The model only compares one row of each chart (heights) before deciding that the vehicle is in compliance, when the weight row of the vehicle’s actual capabilities violates the corresponding weight row of the requirement chart.
When comparing the *-RAG models, we find that, in a departure from other benchmark subsets where GPT-4o-RAG is the best performer, Claude-Opus-RAG performs the best on Functional Performance accuracy, with a score of 0.875. The model also ties GPT-4o-AllRules’s high score on explanation Similarity, further underscoring its good understanding of these questions. This subset of the benchmark required significant technical expertise, perhaps indicating that Claude-Opus has one of the stronger engineering backgrounds of the MLLMs tested. It would be interesting to test and compare Claude-Opus-AllRules to GPT-4o-AllRules, but given the cost constraints of running Claude-Opus-AllRules mentioned previously, this will be left as future work.
Table 4: Analysis of MLLMs’ performances on the Rule Comprehension questions, broken-down by vehicle-component-mention type. Def refers to definition-components, those defined explicitly within the rule document. MultM refers to multi-mention components, those that appear multiple times within the rule document but are not explicitly defined. NoM refers to no-mention components, those that don’t appear verbatim in the rule document. Trends appear in that all models score the highest on definition-components questions within the Definition subset, while all models score the highest on the multi-mention component questions within the Presence subset.
Subset Mention Type Model
GPT-4o- AllRules GPT-4- AllRules GPT-4o- RAG GPT-4- RAG LLaVA- 1.5- RAG Gemini- 1.0- RAG Claude- Opus- RAG
Definition (F1 BoC $\uparrow$) Def 0.78 0.74 0.69 0.56 0.50 0.51 0.48
MultM 0.47 0.42 0.47 0.39 0.37 0.48 0.40
NoM 0.54 0.35 0.54 0.35 0.37 0.48 0.46
Presence (ACC $\uparrow$) Def 0.67 0.58 0.5 0.50 0.42 0.50 0.50
MultM 0.73 0.68 0.78 0.55 0.53 0.63 0.50
NoM 0.80 0.50 0.70 0.50 0.40 0.30 0.50
5 Discussion
In testing various state-of-the-art models (at the time of writing) on DesignQA, we developed questions about what factors impact models’ scores on the benchmark. For example, to what extent is a model’s score on DesignQA influenced by the RAG system used? How does the way in which a component is mentioned in the rule document affect a model’s ability to answer questions about the component? Are models able to better answer questions about engineering drawings when dimensions are shown via scale-bars or when they are directly indicated on the drawing? Does additional image and textual context help models answer questions about rule compliance? We explore these questions in this section, and we propose general suggestions for how to develop models that are better suited to answering the questions presented within DesignQA.
Impact of RAG System on Models’ DesignQA Scores The focus of our work was to develop a benchmark to assess MLLMs’ understandings of technical documentation and to utilize the benchmark to illustrate strengths and weaknesses of state of the art models, at the time of writing. As such, we did not invest time into tuning or developing RAG solutions designed specifically for DesignQA; development of useful RAG methods is a science within itself. [44] However, one of our findings in the previous section was that the simple LlamaIndex RAG used for the *-RAG models often did not provide the MLLM with useful context from the FSAE rule document.
To help disentangle RAG efficacy from inherent model performance, we generated GuaranteedRAG for each question in the benchmark. Created using a series of scripts with key-word matching, GuaranteedRAG is guaranteed to contain context that is relevant to the question at hand. For example, in the case of the Retrieval questions, GuaranteedRAG context is designed to contain the rule-in-question. More details about the implementation of GuaranteedRAG can be found in Appendix B. Due to the cost of running the closed-source models, we only tested the open-source model, LLaVA-1.5, with GuaranteedRAG (LLaVA-1.5-GuaranteedRAG). We generated our GuaranteedRAG five different times (to account for randomness, see Appendix B), and we tested LLaVA-1.5 with each of these five GuaranteedRAGs, reporting averages and standard deviations across the five GuaranteedRAGs.
The results of this experiment can be seen in Table 3. While we expected LLaVA-1.5’s score to improve substantially when tested with GuaranteedRAG instead of simple RAG, LLaVA-1.5-GuaranteedRAG – across most subsets of the benchmark – performs only slightly better than LLaVA-1.5-RAG. When compared with all the other models tested (Table 2), LLaVA-1.5-GuaranteedRAG does not perform the best across any subset of the benchmark. This finding underscores that much of the DesignQA score, at least in the case of LLaVA-1.5, is a result of inherent model capabilities (and lack thereof) rather than RAG efficacy.
The biggest improvement in score between LLaVA-1.5-RAG and LLaVA-1.5-GuaranteedRAG is for the Retrieval questions (0.112 versus 0.699 F1 BoW). Given useful rule document context, the Retrieval questions become straightforward. However, despite receiving the rule-in-question verbatim within an 8716-character text snippet (about 3% of the total rule document), LLaVA-1.5-GuaranteedRAG still struggles to reliably return the rule-in-question for the Retrieval questions. LLaVA-1.5-GuranteedRAG still performs worse on the Retrieval questions than the best performer, GPT-4o-AllRules, which has to search for the relevant rule in the much longer, complete rule text. This finding illustrates LLaVA-1.5’s inferior ability, compared with GPT-4o’s, to extract relevant information from a body of text.
While receiving relevant context helps LLaVA-1.5 perform (on the whole) better on DesignQA, this experiment underscores that high quality RAG may not be the most important factor for improving MLLM scores on the benchmark. LLaVA-1.5 would need significant improvements in inherent model skills – image analysis, engineering knowledge, extraction of relevant information from sections of text – in order to improve its score further on DesignQA. We hope that GuaranteedRAG – which helps isolate the testing of inherent model skills from the testing of RAG implementation – can serve as a useful diagnostic tool for researchers seeking to improve foundation models.
Table 5: Effect of dimensioning system used in engineering drawings on model accuracy on Dimension QAs. These results highlight the fact that the MLLMs tested are generally better able to answer questions about engineering drawings that are direct-dimensioned rather than those that have a scale-bar. An exception is Gemini-1.0-RAG, which performs better on the scale-bar-dimensioned drawings.
ACC by Dimension System Model
GPT-4o- AllRules GPT-4- AllRules GPT-4o- RAG GPT-4- RAG LLaVA- 1.5- RAG Gemini- 1.0- RAG Claude- Opus- RAG
Direct 0.79 0.66 0.66 0.45 0.41 0.51 0.54
Scale-bar 0.75 0.28 0.70 0 0.40 0.55 0.45
Effect of Vehicle Component Mention-Type on the Rule Comprehension Score
We were curious to understand if the manner in which and the number of times a vehicle component is mentioned in the rule document impacts a model’s ability to answer questions about that component. For the Rule Comprehension questions, the component-in-question in each QA was tracked based on how it was mentioned in the rule document (definition-component, multi-mention component, or no-mention component), as explained previously in Section 3.1.3. Table 4 breaks down each model’s score for the Rule Comprehension questions (both Definition and Presence questions) by the three different component-mention types.
There is a clear trend: for the Definition questions, where the model is asked to identify the name of a highlighted vehicle component in an image, each model always scores the highest on the definition components. Recall that for the Definition questions, the *-RAG models did not receive any context from the rule document, since the simple RAG implemented always returned the same section of the rule document (as the text of the prompt was identical for each Definition question). The models’ higher scores on the definition-component questions within the Definition subset might then be explained by the fact that the models have seen similar images of highlighted definition-components during their pre-training. For example, images exist online that highlight the location of vehicle components like the main hoop and the front hoop. Since the definition-components are somewhat specific terms to the Formula SAE competition, they tend to be the focus of the online content and visuals that discuss the Formula SAE rules.
For the Presence questions – where the model is asked to identify (yes/no) whether the specified component is present in an image – the models exhibit a different trend. Almost all models score the highest on the multi-mention components (except for GPT-4o-AllRules and Claude-Opus-RAG, which scores equivalently on all component types). These questions show the models views of the vehicle – zoomed views – for which there are no similar images online. As such, the models are best able to visually identify the multi-mention components, which tend to be more common vehicle terms (like pedal assembly, brake system, and rocker arms).
Regardless of the specific reason for this observed pattern, the results presented in Table 4 reflect that there are trends associated with component-mention type in the rule document and model score on the Rule Comprehension questions. It would be interesting to explore this trend further in future work.
Effect of Engineering Drawing Dimensioning System on Models’ DesignQA Scores
As explained in Section 3.1.2, one-third of the questions comprising the Dimension QAs used scale-bars in the engineering drawings, while the other two-thirds used direct dimensions. We were curious to see what impact these two different dimensioning systems had on model performance. As seen in Table 5, most models perform better on the direct-dimensioned drawings than on the scale-bar drawings. The models seemed to struggle with the scale-bar. For example, GPT-4 sometimes indicated in its explanations that it was using the scale bar for “estimated” dimensions rather than precise ones and in some cases explained that its “image capabilities…do not include measuring dimensions.” GPT-4’s difficulty with image measurement seems to have been improved in the newer version of the model, GPT-4o. GPT-4o-RAG actually had a higher accuracy score for the scale-bar-dimensioned images than for the direct-dimensioned images. Gemini-1.0-RAG was also an outlier, in that it performed better on the scale-bar dimensioned questions than on the direct dimensioned questions. Its ability to answer some of the scale-bar questions correctly likely contributed to the fact that it was the second best performer of the *-RAG models tested. Future research and model development must prioritize this feature, ensuring models can accurately interpret and apply dimensional data across various engineering tasks. This focus will enhance the precision and applicability of MLLMs in complex, real-world engineering environments.
Effect of Additional Context on Models’ Dimension Subset Scores
We explained in Section 3.1.4 that half of the Dimension questions were given additional context that we believed would help in answering the question while the other half were not. This additional context was in the form of highlights in the multiview CAD images – on vehicle components pertaining to the rule-in-question – and corresponding prompt text that revealed the name of the highlighted component (e.g., “the front hoop is highlighted in pink”). Surprisingly, we did not see any obvious trends in model performance with versus without the additional context. GPT-4o-AllRules, GPT-4-AllRules, and LLaVA-1.5-RAG performed worse with additional context, while GPT-4o-AllRules, GPT-4-RAG, Gemini-1.0-RAG, and Claude-Opus-RAG performed better. More investigation into the effect of additional context is needed.
Future Work: Recommendations for Models for This Benchmark
@@ 5 Discussion
In the results presented thus far, we have demonstrated how state-of-the-art models, at the time of writing, perform against our benchmark. Here, we present some observations on how one might modify these models to achieve improved performance on our benchmark.
Firstly, inherent MLLM model skills must be improved in order to achieve higher scores on DesignQA. The models we tested with *-AllRules and *-GuaranteedRAG illustrate that even when given the relevant rule context, MLLMs have skills that need to be strengthened in order to improve their performances on design-according-to-technical-documentation-type tasks. We observed that the models tested struggle to reliably reference rules in the Formula SAE rule document. Rule extraction is a critical skill that models should be able to perform flawlessly; failure to do so could result in catastrophic consequences for subsequent rule compliance questions. Fine-tuning methods, like LoRA [45], QLoRA [46], or InstructLAB [47], could be performed on models using the Retrieval QAs. This process may help MLLMs better understand and “memorize” the rule document, resulting in improved performances across the entire benchmark.
Fine-tuning could also be used to improve MLLMs’ visual recognition of technical components and to strengthen their abilities to analyze engineering drawings. However, in our context, fine-tuning has several limitations. It requires a large amount of data that is not always available, typically struggles with tasks requiring precise recall, and becomes ineffective when design documents are frequently updated. Approaches like retrieval-augmented generation or leveraging expanded context windows of newer models align better with the needs of design-requirement tasks without the constraints associated with fine-tuning.
Second, improvement of the RAG model is critical so that relevant portions of the rule document can be accessed. We saw that GPT-4o-AllRules and GPT-4-AllRules, given all the context, almost always performed better than their respective -RAG models, which were given the selected rule document pages through simple retrieval. While one might argue that these findings illustrate that models with longer context lengths should be developed and used, it is important to note that using models with longer context inputs is significantly more expensive, as of the time of writing. The cost to feed the entire rule document into the context for each question was more than ten times the cost of selecting relevant pages via RAG. As such, effective RAG could help reduce the computational burden and provide useful context to the user regarding relevant sections in the original documents.
While MLLMs coupled with GuarnateedRAG showed improvements over those paired with simple RAG, the GuarnateedRAG solution – which was custom-developed for the questions in DesignQA using key-word matching scripts – is not generalizable to other kinds of design-requirement-type questions. Multimodal RAG might be an interesting avenue for exploration, since the images in the VQAs could help in selecting the relevant portion of the rule document. The RAG could also perhaps be improved by experimenting with different chunking methods. Nevertheless, as models become cheaper, more computationally efficient, and trained with larger context windows, RAG approaches might become less useful.
@@
Engineering design, inherently grounded in practical applications, necessitates a deep understanding of technical documents to ensure that designs not only meet but also adhere to stringent requirements and standards. One of the critical insights emerging from our benchmark evaluations is the realization that modern AI models are still in the early stages of truly understanding engineering documentation. This limitation highlights a significant gap in the path toward fully automated design processes. There is therefore substantial future work to be done to improve MLLM models’ abilities to solve engineering design and design requirement problems.
Limitations
While this benchmark presents a first step into formally evaluating a model’s ability to interpret and understand design requirement documentation, DesignQA does focus on just a single rule document, and it is unknown how the results on DesignQA would vary if a different technical document was used for dataset curation. Unfortunately, benchmark development costs (requiring expert manual QA creation and validation) pose limitations on scaling the size of the dataset or the number of technical documents referenced. Our benchmark is also limited to six types of design-requirement-based tasks, while there are certainly others that are worth exploring. However, we believe that the six subsets of our benchmark are reflective of the primary ways in which engineers consult design requirement documentation. We also inherently made choices about prompt format and CAD/engineering-drawing image presentation in creating our benchmark. DesignQA results may not generalize to similar questions where the prompts or images are tweaked. Exploring the effect of other prompt and image formats on models’ scores on DesignQA could be the subject of future studies. Overall, generalizability (to other similar questions not contained within the test dataset) is a common limitation of LLM benchmarks [48]. While generalizability may be a limitation, DesignQA is an attempt to track MLLMs’ abilities to answer questions about design requirements.
In terms of the validity of DesignQA, the automatic evaluation metrics we used could be viewed as a limitation. There are cases where a model response’s score on an automatic evaluation metric may not match the validity a human would assign to the same response. For example, instances where a model doesn’t follow the specified response format (e.g. not separating Compilation question rules with commas) result in artificially low scores for the model on that metric. There were also cases, for the Retrieval questions, where models returned rules that included child rules (which is not technically wrong), whereas our ground truth reference did not include child rules. Evaluating semantic equivalence is a difficult problem, and the BLEU and ROUGE scores used to evaluate model explanations have known limitations. Collecting more human-generated reference explanations, while time consuming, would allow for more accurate scoring of model explanations. While automated evaluation metrics may slightly artificially penalize models, the alternative – manually scoring responses – would be extremely resource-intensive for thousands of questions and may also have many biases. In fact, automatic evaluation is necessary for a benchmark dataset that is meant to be used by the community to continuously test models developed in the future.
Another limitation of DesignQA that may affect benchmark validity is the small size of certain benchmark subsets. While there are more than a thousand Retrieval questions – since they could be generated automatically – the other subsets had to be manually generated and thoroughly reviewed and are therefore smaller in size. In particular, the Functional Performance questions only have 16 QAs due to the difficulty of obtaining vehicle testing data that was immediately relevant to the requirement document. While some of the benchmark subsets may be small, we prioritized manual, high-quality expert QA creation and validation over alternative methods – like synthetic data generation or crowdsourcing – which may compromise dataset quality.